OpenStudy (handsomechickenlover):

When the distributive property is applied, which number completes the statement correctly?

OpenStudy (handsomechickenlover):

–7 • (–4 + 3.2) = –7 • (–4) ____• 3.2

OpenStudy (handsomechickenlover):

Help! I medal!

OpenStudy (handsomechickenlover):

I need to know what number goes in the ____

Vocaloid (vocaloid):

a*(b+c) = a*b + a*c

OpenStudy (handsomechickenlover):

So the number is 7?

Vocaloid (vocaloid):

not quite, check your sign

OpenStudy (handsomechickenlover):

-7

Vocaloid (vocaloid):

yup, good job

OpenStudy (handsomechickenlover):

Thank you very Much!